# GR-MG: Leveraging Partially Annotated Data via Multi-Modal Goal Conditioned Policy

### v1

@@ III-C Training
For the training of goal image generation model, we follow the training setting in InstructPix2Pix [26] and train a noise prediction model as in DDPM [34]. We sample the frame which is $N$ steps away from the current frame in the trajectory/video as the goal image. As a rough task progress can already be very informative and can bring about more stable training, we discretize the task progress predicted from the policy from 0% to 100% into 10 bins before passing it to the goal image generation model.
For the training of multi-modal goal-conditioned policy, we initialize its weight with the pre-trained model weight derived from the generative video pre-training on Ego4d [25] in alignment with GR-1 [4]. Subsequently, leveraging our multi-modal goals, the policy can be first trained on data without language annotations, e.g., robot play data. Specifically, we input a null string for the text input in this stage. After that, we can further finetune the policy with fully annotated robot trajectories.
The input task progress in the training of the goal image generation model is computed from the timestep of the image in the video. The goal image in the training of the policy is sampled from the ground truth goal image from the trajectory. Therefore, the training of the two modules are independent, although they are interconnected during inference. We leave the investigation of co-training the two modules as future work.
@@ III-D Inference
During inference, the task progress is initially set as zero. And the goal image generation model uses the current observation image, language instruction, and the task progress to generate goal images. The goal image is passed to the multi-modal goal-conditioned policy together with the language instruction, sequence of observation images, and sequence of robot states to predict action trajectories and task progress. As the goal image is set as $N$ steps away from the current image in training, we run the goal image generation model only every $n<N$ steps in a closed-loop manner for efficiency.
@@ IV Experiments
We perform extensive experiments in a simulation benchmark and the real world to evaluate the performance of GR-MG. Throughout the experiments, we aim to answer three questions. 1) Can GR-MG perform multi-task learning and deal with challenging settings including unseen distractors, unseen backgrounds, and unseen objects? 2) Does incorporating robot play data enhance the performance of GR-MG, especially when fully annotated data are scarce? 3) How does the inclusion of videos without action labels improves the performance of GR-MG?
TABLE I: Results On Calvin ABC->D
Method Fully Annotated Data Use Play Data No. Instructions in a Row (1000 chains)
@@
GR-MG w/o lang 100% $\times$ 91.8% 79.8% 68.9% 58.1% 48.1% 3.46±0.04
GR-MG w/o prog 100% $\times$ 94.1% 85.0% 75.2% 65.4% 56.3% 3.76±0.11
@@
GR-MG w/o play data 10% $\times$ 82.4% 60.8% 42.2% 28.7% 19.7% 2.33±0.04
@@
  • "100%" and "10%" denote the proportion of fully annotated data used for training. "w/o prog" indicates that task progress is not provided to the goal image generation model. "w/o lang" indicates that the policy uses only the goal image as the condition, with the language replaced by an empty string. "w/o image" indicates that the policy uses only the language as the condition, with the goal image replaced by a zero image.
Figure 3: Experiments. CALVIN benchmark consists of 34 tasks across four different environments. Real-robot experiments encompass 47 tasks, including pick-and-place operations, manipulation of articulated objects and pouring task.
@@ IV-A1 Experiment Settings
We first perform experiments in the challenging CALVIN [14] benchmark environment. CALVIN focuses on language-conditioned visual robot manipulation. It contains 34 tasks and four different environments, i.e., Env A, B, C, and D (Fig. 3). Each environment contains a Franka Emika Panda robot with a parallel-jaw gripper and a desk for performing different manipulation tasks. Real-robot rollout videos can be found in the supplementary material and the project page.
@@ IV-A2 Generalization to Unseen Environments
We compare GR-MG with various state-of-the-art methods, including one that utilizes 3D signals [36]. The results are presented in Tab. I. Baseline results are taken from the original papers or provided by the authors. GR-MG outperforms all the comparing methods, establishing a new state-of-the-art. The average length reported in the last column of Tab. I serves as a comprehensive metric, indicating the average number of tasks the robot is able to accomplish in a row of five across the evaluated 1,000 sequences. GR-MG improves the average length from 3.35 to 4.04. Additionally, it increases the success rate of completing one and five tasks in a row from 93.8% to 96.8% and from 41.2% to 64.4%, respectively. These results demonstrate GR-MG’s strong multi-task learning capabilities and its ability to generalize to unseen environments.
Figure 4: Visualization of the generated goal images in CALVIN Benchmark. The generated images of GR-MG closely align with the ground truth. However, without the progress information during the generation process, the generated goal images can significantly diverge from the language instruction.
@@ IV-A4 Ablation Studies
[…]
@@ IV-B Real-Robot Experiments
Figure 5: Real-Robot Experiment Settings. We evaluate four settings in the real-robot experiments. 1) Simple: This setting aligns with the training dataset. 2) Unseen Distractor: Unseen distractors are added to the scene. 3) Unseen Background: The background is altered by placing two new tablecloths on the table. 4) Unseen Object: The manipulated objects are replaced with randomly chosen unseen objects. The language instructions are modified correspondingly.
Figure 6: Success Rates of Real-Robot Experiments. On the last column, we show the average success rate of the three unseen generalization settings.
@@ IV-B1 Experiment Settings
To evaluate the performance of GR-MG in the real world, we perform experiments on a real robot platform. The platform consists of a Kinova Gen-3 robot arm equipped with a Robotiq 2F-85 parallel-jaw gripper and two cameras, i.e., one static camera for capturing the workspace and another camera mounted on the end-effector. We collected about 15k human demonstrations across 29 tasks, which include pick-and-place manipulation, articulated object manipulation, and pouring. See Fig. 3 for some examples. We use these data to train GR-MG.
In this experiment, we aim to verify the effectiveness of leveraging partially annotated data without action labels in the training of goal image generation model. Some robot datasets have action labels. But their embodiments and coordinate settings are different from the target platform, making them hard for direct usage in model training. However, if the language labels are provided, they can be used as partially annotated data without action labels. In this experiment, we train the goal image generation model with our real robot data, Something-Something-V2 [23], and the videos extracted from RT-1 dataset [1, 37]. We note that, in general, any video with text annotations can be included in the training.
We design four different settings to evaluate the model performance (see Fig. 5): Simple, Unseen Distractor, Unseen Background, and Unseen Object. In Simple, the scene is set to be similar to those in the training data. In Unseen Distractor, unseen distractors are added to the scene. This setting is challenging because the model may be distracted by these new elements, especially if they have similar colors. In Unseen Background, the background is modified by introducing two new tablecloths that were not present in the training data. In Unseen Object, the robot is instructed to manipulate objects that were not included in the training dataset. And the language instructions are adjusted accordingly, i.e., the language instructions are also unseen. In total, we evaluate 47 different language instructions: 29 of which were seen during training, while the remaining tasks were unseen. We compare the performance of GR-MG with GR-1 [4] and a variant of GR-MG, GR-MG w/o part. ann. data, which uses only the fully annotated data for training the goal image generation model. To ensure a fair comparison, given that GR-1 can also utilize text-annotated videos via generative video pre-training, we incorporate Something-Something-V2 and RT-1 into its pre-training.
@@ IV-B2 Results
Results are shown in Fig. 6. Real-robot rollout videos can be found in the supplementary material and the project page. GR-MG outperforms GR-1 and GR-MG w/o part. ann. data in both simple and all the generalization settings. On average, GR-MG improves the average success rate of GR-1 from 62.5% to 75.0% in Simple setting and 42.4% to 57.6% in the three generalization settings. By comparing with GR-MG w/o part. ann. data, we can conclude that additional text-video data is crucial for enhancing the generalization capability of GR-MG. The failure case of GR-1 generally includes 1) picking wrong objects and 2) closing the gripper too early before arriving at the object. For GR-MG, the failure case mainly results from inaccurate actions, e.g., not able to move into the handle to open the oven in the oven opening task. We believe one reason is that GR-MG does not incorporate any depth information in the model. We plan to investigate incorporating depth information in future work.
@@
To explore how partially annotated text-video data benefits GR-MG, we qualitatively compare the generated goal images between GR-MG and GR-MG w/o part. ann. data (Fig. 7). GR-MG produces high-quality and accurate goal images, while the generated goal images of GR-MG w/o part. ann. data is of comparatively low quality and do not adhere to the language instructions. Specifically, in Simple (Fig. 7 (a)), Unseen Distractor (Fig. 7 (b)), and Unseen Background (Fig. 7 (c)) settings, GR-MG w/o part. ann. data is more likely to generate images that fail to follow language instructions. In the Unseen Object setting (Fig. 7 (d)), the absence of additional partially annotated data increases the likelihood of hallucination: objects from the training set will unexpectedly appear in the generated goal images, even if they are not present in the current scene. These finding indicates that incorporating more text-annotated videos into training enhances 1) the understanding of the semantic in languages and 2) the robustness against out-of-distribution data.
Figure 7: Visualization of Generated Goal Images in Real-Robot Experiments. By utilizing additional text-annotated videos to train the goal image generation model, the generated images exhibit improved visual quality and better align with the language instructions.
@@ V Conclusions
In this paper, we introduce GR-MG, a novel method that utilizes multi-modal goals to predict actions. GR-MG uses both a language and a goal image to condition the action prediction. It requires only the language as input and generate the goal image based on the language via a goal image generation model. This design enables GR-MG to effectively leverage large amounts of partially annotated data that are missing either language or action labels. To enhance the robustness of the generated goal images, we incorporate task progress information in the goal image generation model. GR-MG demonstrates exceptional performance in both simulation and real-world experiments, showing strong generalization across various out-of-distribution settings. Given the capability to utilize partially annotated data, we plan to scale up the training of both the goal image generation model and the policy in the future. Additionally, we plan to investigate integrating depth information to further improve the accuracy of action prediction.

### v2

@@ III-C Training
To train the goal image generation model, we follow the training setting in InstructPix2Pix [31] and train a noise prediction model as in DDPM [39]. We sample the frame which is $N$ steps away from the current frame in the trajectory/video as the goal image. As a rough task progress can already be very informative and can bring about more stable training, we discretize the task progress predicted from the policy from 0% to 100% into 10 bins before passing it to the goal image generation model. As the training of goal image generation model only requires videos with text annotations, we are able to incorporate data w/o action labels in the training alongside the fully-annotated robot trajectories.
For training the multi-modal goal-conditioned policy, we initialize its weight with the pre-trained model weight derived from the generative video pre-training on Ego4d [10] in alignment with GR-1 [4]. The multi-modal goal conditioning allows the policy to utilize partially-annotated data w/o text labels. Specifically, if data w/o text labels is available, we first train the policy with the data, where a null string is provided as the text condition. After that, we finetune the policy on fully-annotated robot trajectories.
The input task progress in the training of the goal image generation model is computed from the timestep of the image frame in the video. The goal image in the training of the policy is sampled from the ground-truth goal image in the trajectory. Therefore, the training of the two modules are independent. We leave the investigation of jointly training the two modules as future work.
@@ III-D Inference
During inference, the task progress is initially set as zero. And the goal image generation model uses the current observation image, text instruction, and task progress to generate the goal image. The goal image is passed to the multi-modal goal-conditioned policy together with the text instruction, sequence of observation images, and sequence of robot states to predict action trajectories and task progress. As the goal image is set as $N$ steps away from the current image in training, we run the goal image generation model only every $n<N$ steps in a closed-loop manner for efficiency.
@@ IV Experiments
We perform extensive experiments in a simulation benchmark and the real world to evaluate the performance of GR-MG. We aim to answer three questions. 1) Can GR-MG perform multi-task learning and deal with challenging generalization settings including unseen distractors, unseen instructions, unseen backgrounds, and unseen objects? 2) Does incorporating data w/o text labels enhance the performance of GR-MG, especially when fully-annotated data is scarce? 3) How does the inclusion of data w/o action labels improve the performance of GR-MG? 4) Can GR-MG learn novel skills in a few-shot setting?
TABLE I: Results On the ABC->D Splits of CAVLIN Benchmark
Method Fully-Annotated Data Partially-Annotated Data No. Instructions in a Row (1000 chains)
@@
GR-MG w/o text 100% $\times$ 91.8% 79.8% 68.9% 58.1% 48.1% 3.46±0.04
GR-MG w/o progress 100% $\times$ 94.1% 85.0% 75.2% 65.4% 56.3% 3.76±0.11
@@
GR-MG w/o part. ann. data 10% $\times$ 82.4% 60.8% 42.2% 28.7% 19.7% 2.33±0.04
@@
  • "100%" and "10%" denote the proportion of fully-annotated data used for training. "w/o progress" indicates that task progress is not provided for the goal image generation model. "w/o text" indicates that the policy uses only the goal image as the condition. "w/o image" indicates that the policy uses only the text as the condition. "w/o part. ann. data" indicates that the policy does not use the trajectories without text labels for training.
Figure 3: Experiments. CALVIN benchmark consists of 34 tasks across four different environments. Real-robot experiments encompass 58 tasks, including pick-and-place and non-pick-and-place manipulations.
@@ IV-A1 Experiment Settings
We perform experiments in the challenging CALVIN [17] benchmark environment. CALVIN focuses on language-conditioned visual robot manipulation. It contains 34 tasks and four different environments, i.e., Env A, B, C, and D (Fig. 3). Each environment contains a Franka Emika Panda robot with a parallel-jaw gripper and a desk for performing different manipulation tasks.
As indicated by [17], only 1% of the collected trajectories were annotated with texts. The remaining trajectories without text labels contain unstructured play data collected by untrained users with no information about the downstream tasks. These trajectories can be used as data w/o text labels for training the multi-modal goal-conditioned policy of GR-MG. We refer the readers to [17] for more details on the dataset.
@@ IV-A2 Generalization to Unseen Environments
We compare GR-MG with various state-of-the-art methods. Results are shown in Tab. I. Results of baseline methods are sourced from the original papers or provided by the authors. GR-MG outperforms all the comparing baseline methods, establishing a new state-of-the-art. The average length reported in the last column of Tab. I serves as a comprehensive metric, indicating the average number of tasks the robot is able to accomplish in a row of five across the evaluated 1,000 sequences. GR-MG improves the average length from 3.35 to 4.04. Additionally, it increases the success rate of completing one and five tasks in a row from 93.8% to 96.8% and from 41.2% to 64.4%, respectively. These results showcase that GR-MG possesses strong capabilities of multi-task learning and generalization to unseen environments.
Figure 4: Visualization of the generated goal images in CALVIN Benchmark and Real-Robot Experiments. (a) The generated images of GR-MG closely align with the ground truths. Without task progress information, the goal images generated by GR-MG w/o progress diverge from the text instructions. (b) Without partially-annotated data, the generated goal images do not adhere to the text instructions and suffer from hallucination.
@@ IV-A4 Ablation Studies
To assess the impact of using multi-modal goals as conditions, we compare with two variants, GR-MG w/o text and GR-MG w/o image, which remove the text condition and goal image condition in the policy during training, respectively. We retrain the two variants to ensure fair comparison. The performance of these two variants is similar. GR-MG significantly outperforms both of them, indicating that both modalities are essential for effective policy conditioning. To investigate the effectiveness of task progress, we also compare with GR-MG w/o progress, which excludes the task progress information in the goal image generation model. GR-MG surpasses this variant, showcasing the effectiveness of incorporating task progress information.
In addition, we perform quantitative comparison on the goal images generated by GR-MG and GR-MG w/o progress. We compare the similarity between generated goal images and ground truths using four metrics: Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Cosine Distance of ResNet50 Features (CD-ResNet50). Results are shown in Tab. II. GR-MG outperforms GR-MG w/o progress in all four metrics, underscoring the effectiveness of incorporating task progress information in goal image generation. Qualitative results are shown in Fig. 4(a). Although GR-MG w/o progress is able to generate goal images with high visual quality, these images do not align well with the text instructions. In contrast, GR-MG accurately follows the text instructions and generates goal images that closely match the ground truths.
TABLE II: Ablation Studies on Task Progress for Goal Image Generation
Method MSE ↓ PSNR ↑ SSIM ↑ CD-ResNet50 ↑
GR-MG w/o progress 965.347 18.821 0.721 0.945
GR-MG (Ours) 903.139 19.121 0.730 0.946
@@ IV-B Real-Robot Experiments
Figure 5: Success Rates of Real-Robot Experiments. Unseen Average shows the average success rate of the four unseen generalization settings.
@@ IV-B1 Experiment Settings
To evaluate the performance of GR-MG in the real world, we perform experiments on a real robot platform. The platform consists of a Kinova Gen-3 robot arm equipped with a Robotiq 2F-85 parallel-jaw gripper and two cameras, i.e., one static camera for capturing the workspace and another camera mounted on the end-effector. The training dataset consists of 18k human demonstrations across 37 tasks, which include 23 pick-and-place tasks and 14 non pick-and-place tasks such as pouring, flipping, and rotating. See Fig. 3 for some examples.
In this experiment, we aim to verify the effectiveness of leveraging data w/o action labels in training. Specifically, we combine the data from Something-Something-V2 [8] and RT-1 [1] with our real robot data to create the training dataset for training the goal image generation model. It is worth noting that any videos with text annotations can be included in the training process.
We design five different settings to evaluate the model performance: Simple, Unseen Distractors, Unseen Instructions, Unseen Backgrounds, and Unseen Objects. In Simple, the scene is set to be similar to those in the training data. In Unseen Distractors, unseen distractors are added to the scene. In Unseen Instructions, we follow [5] and use GPT-4 to generate unseen synonyms for the verbs in the instructions. For example, we replace "pick up" with "take", "cap" with "cover", and "stack" with "pile". In Unseen Backgrounds, the background is modified by introducing two new tablecloths that were not present in the training data. In Unseen Objects, the robot is instructed to manipulate objects that were not included in the training dataset. And the language instructions are adjusted accordingly, i.e., the language instructions are also unseen. In total, we evaluate 58 different tasks: 37 of which were seen during training, while the rest were unseen. See the appendix on the project page for the full list of training tasks and the 58 evaluated tasks. We compare the performance of GR-MG with OpenVLA [42], Octo [3], and GR-1 [4]. To study the effectiveness of leveraging data w/o action labels, we also compare with the variant which only uses fully-annotated data for training the goal image generation model, GR-MG w/o part. ann. data.
@@ IV-B2 Results
Results are shown in Fig. 5. Real-robot rollout videos can be found in the supplementary material and on the project page. GR-MG outperforms all the baseline methods in all five evaluation settings. A typical failure mode of OpenVLA is inaccurate grasping. We believe this is because it uses an discretized action space, making the action prediction inaccurate. Also, it does not support inputs of observation history and wrist camera observation, making it difficult to accurately identify the appropriate timestep to close/open the gripper in some tasks. Octo supports inputs of observation history, proprioceptive data, and observation from the wrist camera. However, it is not able to generalize well in unseen settings, especially for Unseen Backgrounds and Unseen Objects. GR-1 performs well in most settings, but it is not able to pick the correct object in the setting of Unseen Objects. GR-MG is able to generalize well to unseen settings. The performance degradation in Unseen Instructions is modest for GR-MG compared to other baseline methods. We believe the powerful generalization capabilities of GR-MG stem from the video pre-training for both the goal image generation model and the policy. It allows the model to harness the prior knowledge gained from pre-training to enhance its performance in downstream manipulation tasks.
@@
By comparing with GR-MG w/o part. ann. data, we conclude that additional partially-annotated data w/o action labels are crucial for enhancing the generalization capabilities of GR-MG. We qualitatively compare the goal images generated by GR-MG and GR-MG w/o part. ann. data in Fig. 4(b). GR-MG generates high-quality and accurate goal images, while the goal images generated by GR-MG w/o part. ann. data do not adhere to the language instructions. In the setting of Unseen Objects (second row of Fig. 4(b)), the absence of additional partially-annotated training data increases the likelihood of hallucination: objects from the training set will unexpectedly appear in the generated goal images even if they are not present in the current scene. These findings indicate that incorporating more data w/o action labels into training enhances 1) understanding of language semantics and 2) robustness against out-of-distribution data.
IV-B3 Few-Shot Learning of Novel Skills
We further perform experiments to evaluate the few-shot learning capabilities of GR-MG. In particular, we first hold out 8 tasks from the 37 tasks and train the model on the rest 29 tasks. The total number of trajectories used for training is 15k. These 8 tasks include 7 novel skills which are unseen in the 29 training tasks. For example, we hold out "wipe the cutting board" and there are no wiping tasks in the 29 training tasks. For the full list of all the held-out tasks, please see the appendix on the project page. After training, we further finetune the model with data containing 10 (or 30) trajectories per held-out task. That is, the total number of trajectories used for finetuning in this stage is 80 (or 240). We evaluate on the 8 held-out tasks and results are shown in Tab. III. GR-MG outperforms all the comparing baseline methods in both 10-shot and 30-shot settings. We observe that the main bottleneck of GR-MG lies within the policy. The goal image generation model is able to generate accurate goal images on the these novel skills after few-shot finetuning, but the policy struggles to perform well. To tackle this challenge, we plan to investigate scaling up the training of the policy with more real-world data w/o text labels in the future.
TABLE III: Success Rates of Few-Shot Learning
      Methods       10-shot       30-shot
      OpenVLA [42]       0.0%       2.5%
      Octo [3]       0.0%       0.0%
      GR-1 [4]       2.5%       22.5%
      GR-MG w/o part. ann. data       10.0%       27.5%
      GR-MG (Ours)       17.5%       37.5%
@@ V Conclusions
In this paper, we introduce GR-MG, a novel method that leverages multi-modal goals to predict actions. GR-MG uses both a language and a goal image to condition the action prediction. It requires only the language as input and generate the goal image based on the language via a goal image generation model. This design enables GR-MG to effectively leverage large amounts of partially-annotated data that are missing either text or action labels. To improve the robustness of the generated goal images, we incorporate task progress information in the goal image generation model. GR-MG showcases exceptional performance in both simulation and real-world experiments, showing strong generalization capabilities across various out-of-distribution settings. It can also efficiently learn novel skills in a few-shot setting. We plan to scale up the training of both the goal image generation model and the policy by incorporating more partially-annotated data in the future. Additionally, we plan to investigate integrating depth information to further improve the accuracy of action prediction.
VI ACKNOWLEDGEMENT
This work was partially supported by National Natural Science Foundation of China (62322607, 62236010 and 62276261), and Youth Innovation Promotion Association of Chinese Academy of Sciences under Grant 2021128.